OpenStudy (anonymous):

Find the value of 5 · 23. 30 40 1,000

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

? That isn't an answer choice

OpenStudy (anonymous):

I believe this is a troll Q.. you perfectly know how to multiply. (Refer to your prior posts(s))

OpenStudy (mathstudent55):

\(5 \times 2^3\) Do the 2^3 part first. 2^3 = 2 * 2 * 2

OpenStudy (anonymous):

i meant the third answer

OpenStudy (anonymous):

I thought the Q was 5 times 23. My apologies.

OpenStudy (mathstudent55):

It is not the third answer. You need to do 2^3 first. 2^3 = 2 * 2 * 2 = 8 Then you do 5 * 8 What is 5 * 8? That is the answer.